OpenStudy (anonymous):

Look at the relationship between a and b. a 1 3 5 7 b 52 50 48 46 Which equation below describes the relationship between a and b? A. b = 52 × a B. b = 52 ÷ a C. b = 51 + a D. b = 53 - a

OpenStudy (anonymous):

D.

OpenStudy (welshfella):

just plug in values for b and a and see which fit for example try A; b = 52 * a try the first values of a and b 52 = 52 * 1 = 52 - this fits second set of values 50 = 52 * 3 = 156 - doesn't fit so we know the answer is not A

OpenStudy (welshfella):

@sourav_aich - plz don't just give the answers - jayonna15 learns nothing from this!!